# Observable Primordial Gravitational Waves from Non-minimally Coupled $R^{2}$ Palatini Modified Gravity

### v1

@@ 5 Gravitational Wave Observations
PLICs are detector sensitivity curves for stochastic gravitational-wave backgrounds that take into account the increase in sensitivity that comes from integrating over frequency in addition to integrating over time [116]. The plots in this section show the PLIC for different detectors as shaded backgrounds and the spectrum from primordial gravitational waves as solid lines. The plots can be interpreted as follows: If a curve for a predicted GW background lies below the PLIC, then for such a background, the sound-to-noise ratio (SNR) for that particular detector is $<1$. While, if it lies anywhere above the PLIC, then the background will be observed in the detector with SNR $>1$. PLIC curves for LIGO (O3) and LIGO (O5) are generated using the formalism mentioned in [116] while for BBO, CE, ET, DECIGO, LISA, PTAs, SKA, and NANOGrav is obtained from [117]. The PLICs for ASTROD-GW and $\mu$Ares were obtained from their respective noise spectra [89, 90].
A study considered a model with non-minimal coupling of inflaton and gravity and minimally coupled higher order curvature term as $\mathcal{L}\propto h(\chi)R+\alpha R^{2}$, where $h(\chi)=1+\chi^{2}$ [99]. They claimed that having a longer period of hyperkination boosts the GW spectrum to an observable regime. However, from the energy density argument, the length of hyperkination is bounded from above, and thus, considering a longer period will destabilize the universe. We probe the observability of GWs from their model but with $\Delta N_{\text{hyp}}=0.2$. The left panel in the Figure 7 shows the observability considering $\Delta N_{\text{hyp}}=15$, while the right panel takes $\Delta N_{\text{hyp}}=0.2$. A difference is apparent from the figure, a shorter period of hyperkination gives longer kination and longer period for reheating. While the order of magnitude of the boost is the same, the spectrum gets enhanced in a different frequency range, making it less detectable by future GW experiments.
@@
One of the earliest models of inflation was the Starobinsky model [7], $\mathcal{L}\propto R+\alpha R^{2}$, which we can retrieve from our model by setting $t=0$. The observability of this accepted model is also tested in our formalism. The freedom over choosing $t$ allows us to probe a wider class of theories as high as $\chi^{2}R^{2}$ and their probable detection in the future. We probe the detectability for a few different parameter values for $H$ and $\Omega_{r}^{\text{end}}$ from the available parameter space in Figure 11. We can see an array of effects from this variation. All cases result in giving a boosted spectrum with a plateaued hyperkination phase. The plateau during reheating also gains energy as we decrease the order of the coupling. The plateau during reheating and the boost during kination fall under the detectability of a number of future GW experiments. The observation of the plateau during reheating will constrain the values of $\Omega_{r}^{\text{end}}$ and $H$ if such is observed in the near future. However, to confirm a particular theory, the observation of boost during kination is required, which will give the spectral shape of the enhancement with respect to frequency. However, the plateau from hyperkination is outside any observable region of current-future detectors. This prompts a need for GW detectors in the kHz to GHz frequency range, which can confirm the presence of the boost from kination and the bound from hyperkination. One proposal to detect ultra-high frequency gravitational waves (UHF-GWs), using inverse Gertsenshtein effect [118], has gained popularity recently. This has placed resonant cavities (RC) as promising candidates for detecting gravitational wave signals from these epochs within the MHz - GHz UHF-GW band [93].
@@ 6 Summary and Conclusions
In this paper, we focused on the possibility of observing primordial gravitational waves produced during the eras of hyperkination, kination, and reheating. We considered a model with a higher order curvature term coupled to the inflaton field of the form $h(\chi)=(1+\xi\chi/M_{\text{Pl}})^{t}$, where we can adjust the order of the coupling $t$. This model features a non-oscillatory inflaton potential that allows the inflaton to roll freely after it comes down from the inflationary plateau. The presence of $R^{2}$ terms in the Einstein frame results in non-canonical kinetic terms that are responsible for hyperkination and the flat and low potential demands movement of the scalar field under kinetic domination from canonical kinetic terms, giving the epoch of kination. We probe a wide class of theories from lower-order $(t<0)$, Starobinsky ($t=0$), and non-minimal coupling between the inflaton field and $R^{2}$ ($t>0$). We formulated the dynamics of the field during those times and obtained a bound on the length of the phase of hyperkination, $\Delta N_{\text{hyp}}\leq 0.2$, and the order of coupling, $t\leq 2$, from the requirement of the field derivative to be a kinetic attractor which restricts up to $\chi^{2}R^{2}$. However, such a model for $t>0$ does not naturally give out reheating. We fixed this issue by providing a supplementary mechanism that governs the decay of the scalar field and brings out the conditions necessary for radiation domination and, thus, reheating. We assumed that the history of the universe is as follows: inflation, hyperkination, kination, reheating (BBN) followed by radiation and matter domination under $\Lambda$CDM.
@@
We then solved the evolution of the scale factor with time during this history. The change in scale factor due to these new epochs brings out changes in tensor mode perturbations and, thus, in the observed GW spectrum. We solved the Mukhanov-Sasaki equation to obtain the perturbations and observed that as soon as the modes enter the kination phase, they obtain a large amount of energy. We also observed that coupling of lesser order boosts the perturbations during hyperkination and kination and, having a smaller duration of hyperkination, dilutes the contribution of GWs from kination. Following the mode calculation, we examined the spectral energy density coming from those primordial GWs. The spectrum features two plateaus, one after inflation during hyperkination and the other during reheating. The two plateaus are connected by the boost, $\Omega_{\text{GW}}\propto f^{2/p}$, gained during the phase of kination. The enhancement changes as a function of the coupling $t$. The epoch of hyperkination prevents the boost in the spectrum from violating the BBN bounds, thus demanding the presence of $R^{2}$ terms in the Einstein-Hilbert action whenever a runaway-like, non-oscillatory potential is considered. We probed the available parameter space for $\alpha$, $H$, and $\Omega_{r}^{\text{end}}$ for different orders of coupling, which are required to check the validity of the theory. Finally, we checked the observability of these modified spectra by current and future GW detectors by superimposing the $\Omega_{\text{GW}}$ spectra against the power-law integrated sensitivity curves. We find that as we decrease the order of the coupling, the spectra shift towards a more observable regime. The observation of the plateau during reheating will constrain the $H$ and $\Omega_{r}^{\text{end}}$ values, while the spectral shape of the boost obtained during kination will confirm the nature of the theory. To probe the bound placed by hyperkination that prevents the boost from kination from violating the Big Bang Nucleosynthesis, GW detector experiments, like detection via RC in the kHz-GHz frequency regime, are needed. The flat spectrum from reheating for lower-order coupling ($t<0$) can be matched with PTA and NANOGrav observations of the SGWB to constrain the parameter space. We can also place bounds on the parameter space detectable by future experiments; we leave these crucial analyses for the future.
@@ Appendix A Derivation for $\bar{\bar{\chi}}$
Appendix B Derivation for Scale Factors
@@
$a(\eta_{\text{end}})=\left[\frac{-1}{(1-\epsilon)H\eta_{\text{end}}}\right]^{1/(1-\epsilon)}=1\implies H\eta_{\text{end}}=-1\because\epsilon=0$ (B.1)
@@
$\rho=\frac{1-e^{2(N-\Delta N_{\text{hyp}})}}{6\alpha\kappa h~{}e^{4(N-\Delta N_{\text{hyp}})}}=\frac{1-a^{2}e^{-2\Delta N_{\text{hyp}}}}{6\alpha\kappa h~{}a^{4}e^{-4\Delta N_{\text{hyp}}}}$ (B.2)
@@
$\rho_{\text{end}}=\frac{1-e^{-2\Delta N_{\text{hyp}}}}{6\alpha\kappa h~{}e^{-4\Delta N_{\text{hyp}}}}=3H^{2}\implies\frac{1-e^{-2\Delta N_{\text{hyp}}}}{18\kappa hH^{2}~{}e^{-4\Delta N_{\text{hyp}}}}=\alpha$ (B.3)
$\mathrm{d}\eta$ $=\mathrm{d}a\sqrt{\frac{18\alpha\kappa h~{}e^{-4\Delta N_{\text{hyp}}}}{1-a^{2}e^{-2\Delta N_{\text{hyp}}}}}$ (B.4)
$\implies\eta$ $=\frac{\sqrt{e^{2\Delta N_{\text{hyp}}}-1}\arcsin\left(e^{-\Delta N_{\text{hyp}}}a\right)}{H}+C$ (B.5)
@@
$C=\eta_{\text{end}}-\frac{\sqrt{e^{2\Delta N_{\text{hyp}}}-1}\arcsin\left(e^{-\Delta N_{\text{hyp}}}\right)}{H}$ (B.6)
@@
$a(\eta)$ $=e^{\Delta N_{\text{hyp}}}\sin{\left(\frac{H\eta+1}{\sqrt{e^{2\Delta N_{\text{hyp}}}-1}}+\arcsin\left(e^{-\Delta N_{\text{hyp}}}\right)\right)}$ (B.7)
@@
$\rho=\frac{c}{3\alpha\kappa h}a^{t\sqrt{6}-6}a^{-\Gamma\tau}=\frac{6cH^{2}~{}e^{-4\Delta N_{\text{hyp}}}}{1-e^{-2\Delta N_{\text{hyp}}}}\cdot a^{t\sqrt{6}-6-\Gamma\tau}\equiv Ba^{t\sqrt{6}-6-\Gamma\tau}$ (B.8)
$\mathrm{d}\eta$ $=\frac{\mathrm{d}a}{a^{2}}\sqrt{\frac{3}{Ba^{t\sqrt{6}-6-\Gamma\tau}}}\implies\eta=\frac{2\sqrt{3}}{\sqrt{B}\left(\Gamma\tau-t\sqrt{6}+4\right)a^{\frac{t\sqrt{6}-4-\Gamma\tau}{2}}}+C$ (B.9)
@@
$C=\eta_{\text{kin}}-\frac{a_{\text{kin}}^{p}\sqrt{3}}{\sqrt{B}p}$ (B.10)
@@
$a_{\text{kin}}\equiv a(\eta_{\text{kin}})=e^{\Delta N_{\text{hyp}}}\sin{\left(\frac{H\eta_{\text{kin}}+1}{\sqrt{e^{2\Delta N_{\text{hyp}}}-1}}+\arcsin\left(e^{-\Delta N_{\text{hyp}}}\right)\right)}$ (B.11)
@@
$a(\eta)=\left(p\frac{He^{-\Delta N_{\text{hyp}}}\sqrt{2c}}{\sqrt{e^{2\Delta N_{\text{hyp}}}-1}}(\eta-\eta_{\text{kin}})+a_{\text{kin}}^{p}\right)^{1/p}$ (B.12)
@@
$\rho=3H^{2}a^{-4}$ (B.13)
$\mathrm{d}\eta=\frac{\mathrm{d}a}{a^{2}}\sqrt{\frac{3}{3H^{2}a^{-4}}}=\frac{\mathrm{d}a}{H}\implies\eta=\frac{a}{H}+C$ (B.14)
@@
$C=\eta_{\text{reh}}-\frac{a_{\text{reh}}}{H}$ (B.15)
@@
$a_{\text{reh}}\equiv a(\eta_{\text{reh}})=\left(p\frac{He^{-\Delta N_{\text{hyp}}}\sqrt{2c}}{\sqrt{e^{2\Delta N_{\text{hyp}}}-1}}(\eta_{\text{reh}}-\eta_{\text{kin}})+a_{\text{kin}}^{p}\right)^{1/p}$ (B.16)
@@
$a(\eta)=H(\eta-\eta_{\text{reh}})+a_{\text{reh}}$ (B.17)
Appendix C Solving for Mode Functions

### v2

@@ 5 Gravitational Wave Observations
PLICs are detector sensitivity curves for stochastic gravitational-wave backgrounds that take into account the increase in sensitivity that comes from integrating over frequency in addition to integrating over time [116]. The plots in this section show the PLIC for different detectors as shaded backgrounds and the spectrum from primordial gravitational waves as solid lines. The plots can be interpreted as follows: If a curve for a predicted GW background lies below the PLIC, then for such a background, the sound-to-noise ratio (SNR) for that particular detector is $<1$. While, if it lies anywhere above the PLIC, then the background will be observed in the detector with SNR $>1$. PLIC curves for LIGO (O3) and LIGO (O5) are generated using the formalism mentioned in [116]; curves BBO, CE, ET, DECIGO, LISA, PTAs, SKA, and NANOGrav is obtained from [117] while the one for Resonant Cavities is from [118]. The PLICs for ASTROD-GW and $\mu$Ares were obtained from their respective noise spectra [84, 85].
5.1 Minimal $R^{2}$ with $\Delta N_{\text{hyp}}=0.2$
A study considered a model with $\mathcal{L}\propto h(\chi)R+\alpha R^{2}$, where $h(\chi)=1+\chi^{2}$ [95]. They claimed that having a longer period of hyperkination boosts the GW spectrum to an observable regime. However, from the energy density argument, the length of hyperkination is bounded from above, and thus, considering a longer period will destabilize the universe. We probe the observability of GWs from their model but with $\Delta N_{\text{hyp}}=0.2$. The left panel in the Figure 7 shows the observability considering $\Delta N_{\text{hyp}}=15$, while the right panel takes $\Delta N_{\text{hyp}}=0.2$. A difference is apparent from the figure, a shorter period of hyperkination gives longer kination and longer period for reheating. While the order of magnitude of the boost is the same, the spectrum gets enhanced in a different frequency range, making it less detectable by future GW experiments.
@@ 5.3 Non-minimal $R^{2}$
The freedom over choosing $t$ allows us to probe a wider class of theories as high as $\chi^{2}R^{2}$ and their probable detection in the future. We probe the detectability for a few different parameter values for $H$ and $\Omega_{r}^{\text{end}}$ from the available parameter space in Figure 11. We can see an array of effects from this variation. All cases result in giving a boosted spectrum with a plateaued hyperkination phase. The plateau during reheating also gains energy as we decrease the order of the coupling. The plateau during reheating and the boost during kination fall under the detectability of a number of future GW experiments.
5.4 Minimal Starobinsky Model
One of the earliest models of inflation was the Starobinsky model [7], $\mathcal{L}\propto R+\alpha R^{2}$, which we can retrieve from our model by setting $t=0$. The observability of this accepted theory is also tested in this work. The centre-left panel in Figure 11 shows the GW observability for the coupling $t=0$ or the minimal Starobinsky model in Palatini formalism. The bottom-left panel in Figure 9 shows parameter space for which the theory remains valid. The availability of a large parameter space, the overlap of the GW spectra with observable region and a wide acceptance of this theory makes it exciting to anticipate future observations.
The observation of the plateau during reheating will constrain the values of $\Omega_{r}^{\text{end}}$ and $H$ if such is observed in the near future. However, to confirm a particular theory, the observation of boost during kination is required, which will give the spectral shape of the enhancement with respect to frequency. However, the plateau from hyperkination is outside any observable region of current-future detectors. This prompts a need for GW detectors in the kHz to GHz frequency range, which can confirm the presence of the boost from kination and the bound from hyperkination. One proposal to detect ultra-high frequency gravitational waves (UHF-GWs), using inverse Gertsenshtein effect [119], has gained popularity recently. This has placed resonant cavities (RC) as promising candidates for detecting gravitational wave signals from these epochs within the MHz - GHz UHF-GW band [88].
@@ 6 Summary and Conclusions
In this paper, we focused on the possibility of observing primordial gravitational waves produced during the eras of hyperkination, kination, and reheating. We considered a model with a higher order curvature term coupled to the inflaton field of the form $h(\chi)=(1+\xi\chi/M_{\text{Pl}})^{t}$, where we can adjust the order of the coupling $t$. This model features a non-oscillatory inflaton potential that allows the inflaton to roll freely after it comes down from the inflationary plateau. The presence of $R^{2}$ terms in the Einstein frame results in non-canonical kinetic terms that are responsible for hyperkination and the flat and low potential demands movement of the scalar field under kinetic domination from canonical kinetic terms, giving the epoch of kination. We probe a wide class of theories from lower-order $(t<0)$, Starobinsky ($t=0$), and non-minimal coupling between the inflaton field and $R^{2}$ ($t>0$). We formulated the dynamics of the field during those times and obtained a bound on the length of the phase of hyperkination, $\Delta N_{\text{hyp}}\leq 0.2$, and the order of coupling, $t\leq 2$, from the requirement of the field derivative to be a kinetic attractor which restricts up to $\chi^{2}R^{2}$. However, such a model for $t>0$ does not naturally give out reheating. We fixed this issue by providing a supplementary mechanism that governs the decay of the scalar field and brings out the conditions necessary for radiation domination and, thus, reheating with $T_{\text{reh}}\sim 10^{8}$ GeV. We assumed that the history of the universe is as follows: inflation, hyperkination, kination, reheating (BBN) followed by radiation and matter domination under $\Lambda$CDM.
@@
We then solved the evolution of the scale factor with time during this history. The change in scale factor due to these new epochs brings out changes in tensor mode perturbations and, thus, in the observed GW spectrum. We solved the Mukhanov-Sasaki equation to obtain the perturbations and observed that as soon as the modes enter the kination phase, they obtain a large amount of energy. We also observed that coupling of lesser order boosts the perturbations during hyperkination and kination and, having a smaller duration of hyperkination, dilutes the contribution of GWs from kination. Following the mode calculation, we examined the spectral energy density coming from those primordial GWs. The spectrum features two plateaus, one after inflation during hyperkination and the other during reheating. The two plateaus are connected by the boost, $\Omega_{\text{GW}}\propto f^{2/p}$, gained during the phase of kination. The enhancement changes as a function of the coupling $t$. The epoch of hyperkination prevents the boost in the spectrum from violating the BBN bounds, thus demanding the presence of $R^{2}$ terms in the Einstein-Hilbert action whenever a runaway-like, non-oscillatory potential is considered. We probed the available parameter space for $\alpha$, $H$, and $\Omega_{r}^{\text{end}}$ for different orders of coupling, which are required to check the validity of the theory. Finally, we checked the observability of these modified spectra by current and future GW detectors by superimposing the $\Omega_{\text{GW}}$ spectra against the power-law integrated sensitivity curves. We find that as we decrease the order of the coupling, the spectra shift towards a more observable regime. We also highlighted the future observability of GWs from minimal Starobinsky model in Palatini formalism. The observation of the plateau during reheating will constrain the $H$ and $\Omega_{r}^{\text{end}}$ values, while the spectral shape of the boost obtained during kination will confirm the nature of the theory. To probe the bound placed by hyperkination that prevents the boost from kination from violating the Big Bang Nucleosynthesis, GW detector experiments, like detection via RC in the MHz-GHz frequency regime, are needed. The flat spectrum from reheating for lower-order coupling ($t<0$) can be matched with PTA and NANOGrav observations of the SGWB to constrain the parameter space. We can also place bounds on the parameter space detectable by future experiments; we leave these crucial analyses for the future.
@@ Appendix A Derivation for $\bar{\bar{\chi}}$
Appendix B $\chi^{2}R^{2}$ Inflation
For, $V=\beta\chi^{2}$ and $h=\gamma\chi^{2}$ we obtain the following parameters for this case,
$\epsilon_{\text{v}}=\frac{2}{\chi^{2}}\hskip 8.5359pt\text{and}\hskip 8.5359pt|\eta_{\text{v}}|=\frac{6}{\chi^{2}}$ (B.1)
$n_{s}=1-\frac{24}{\chi^{2}}\hskip 8.5359pt\text{and}\hskip 8.5359ptr=\frac{32}{\chi^{2}}$ (B.2)
$U(\chi)=\frac{\beta}{\gamma\chi^{2}(8\alpha\kappa\beta+\gamma)}\hskip 8.5359pt\text{and}\hskip 8.5359pt\chi(\chi)=\frac{\ln{\chi}}{\sqrt{\gamma+8\alpha\kappa\beta}}$ (B.3)
Figure 12: Tensor-to-scalar ratio $r$ plotted against the primordial tilt $n_{s}$ with Planck observations [18] for $V\propto\chi^{2}$ and $h\propto\chi^{2}$.
Figure 12 shows the Tensor-to-scalar ratio $r$ plotted against the primordial tilt $n_{s}$ with Planck observations for the considered case. Thus we find this model as valid and claim that successful inflation is observed in $\chi^{2}R^{2}$ theory with a quadratic potential.
Appendix C Derivation for Scale Factors
@@
$a(\eta_{\text{end}})=\left[\frac{-1}{(1-\epsilon)H\eta_{\text{end}}}\right]^{1/(1-\epsilon)}=1\implies H\eta_{\text{end}}=-1\because\epsilon=0$ (C.1)
@@
$\rho=\frac{1-e^{2(N-\Delta N_{\text{hyp}})}}{6\alpha\kappa h~{}e^{4(N-\Delta N_{\text{hyp}})}}=\frac{1-a^{2}e^{-2\Delta N_{\text{hyp}}}}{6\alpha\kappa h~{}a^{4}e^{-4\Delta N_{\text{hyp}}}}$ (C.2)
@@
$\rho_{\text{end}}=\frac{1-e^{-2\Delta N_{\text{hyp}}}}{6\alpha\kappa h~{}e^{-4\Delta N_{\text{hyp}}}}=3H^{2}\implies\frac{1-e^{-2\Delta N_{\text{hyp}}}}{18\kappa hH^{2}~{}e^{-4\Delta N_{\text{hyp}}}}=\alpha$ (C.3)
$\mathrm{d}\eta$ $=\mathrm{d}a\sqrt{\frac{18\alpha\kappa h~{}e^{-4\Delta N_{\text{hyp}}}}{1-a^{2}e^{-2\Delta N_{\text{hyp}}}}}$ (C.4)
$\implies\eta$ $=\frac{\sqrt{e^{2\Delta N_{\text{hyp}}}-1}\arcsin\left(e^{-\Delta N_{\text{hyp}}}a\right)}{H}+C$ (C.5)
@@
$C=\eta_{\text{end}}-\frac{\sqrt{e^{2\Delta N_{\text{hyp}}}-1}\arcsin\left(e^{-\Delta N_{\text{hyp}}}\right)}{H}$ (C.6)
@@
$a(\eta)$ $=e^{\Delta N_{\text{hyp}}}\sin{\left(\frac{H\eta+1}{\sqrt{e^{2\Delta N_{\text{hyp}}}-1}}+\arcsin\left(e^{-\Delta N_{\text{hyp}}}\right)\right)}$ (C.7)
@@
$\rho=\frac{c}{3\alpha\kappa h}a^{t\sqrt{6}-6}a^{-\Gamma\tau}=\frac{6cH^{2}~{}e^{-4\Delta N_{\text{hyp}}}}{1-e^{-2\Delta N_{\text{hyp}}}}\cdot a^{t\sqrt{6}-6-\Gamma\tau}\equiv Ba^{t\sqrt{6}-6-\Gamma\tau}$ (C.8)
$\mathrm{d}\eta$ $=\frac{\mathrm{d}a}{a^{2}}\sqrt{\frac{3}{Ba^{t\sqrt{6}-6-\Gamma\tau}}}\implies\eta=\frac{2\sqrt{3}}{\sqrt{B}\left(\Gamma\tau-t\sqrt{6}+4\right)a^{\frac{t\sqrt{6}-4-\Gamma\tau}{2}}}+C$ (C.9)
@@
$C=\eta_{\text{kin}}-\frac{a_{\text{kin}}^{p}\sqrt{3}}{\sqrt{B}p}$ (C.10)
@@
$a_{\text{kin}}\equiv a(\eta_{\text{kin}})=e^{\Delta N_{\text{hyp}}}\sin{\left(\frac{H\eta_{\text{kin}}+1}{\sqrt{e^{2\Delta N_{\text{hyp}}}-1}}+\arcsin\left(e^{-\Delta N_{\text{hyp}}}\right)\right)}$ (C.11)
@@
$a(\eta)=\left(p\frac{He^{-\Delta N_{\text{hyp}}}\sqrt{2c}}{\sqrt{e^{2\Delta N_{\text{hyp}}}-1}}(\eta-\eta_{\text{kin}})+a_{\text{kin}}^{p}\right)^{1/p}$ (C.12)
@@
$\rho=3H^{2}a^{-4}$ (C.13)
$\mathrm{d}\eta=\frac{\mathrm{d}a}{a^{2}}\sqrt{\frac{3}{3H^{2}a^{-4}}}=\frac{\mathrm{d}a}{H}\implies\eta=\frac{a}{H}+C$ (C.14)
@@
$C=\eta_{\text{reh}}-\frac{a_{\text{reh}}}{H}$ (C.15)
@@
$a_{\text{reh}}\equiv a(\eta_{\text{reh}})=\left(p\frac{He^{-\Delta N_{\text{hyp}}}\sqrt{2c}}{\sqrt{e^{2\Delta N_{\text{hyp}}}-1}}(\eta_{\text{reh}}-\eta_{\text{kin}})+a_{\text{kin}}^{p}\right)^{1/p}$ (C.16)
@@
$a(\eta)=H(\eta-\eta_{\text{reh}})+a_{\text{reh}}$ (C.17)
Appendix D Solving for Mode Functions